-The equilibrium constant, Kc, for the following reaction is 5.10×10-6 at 548 K.

NH4Cl(s) NH3(g) + HCl(g)

Calculate the equilibrium concentration of HCl when 0.452 moles of NH4Cl(s) are introduced into a 1.00 L vessel at 548 K.

[HCl] = _________M

-The equilibrium constant, Kc, for the following reaction is 55.6 at 698 K.

H2(g) + I2(g)  2 HI (g)

Calculate the equilibrium concentrations of reactants and product when 0.380 moles of H2and 0.380 moles of I2are introduced into a 1.00 L vessel at 698 K.

[ H2] = _______M
[ I2] = _______M
[ HI ] = _______M

THANK YOU
Answer #1

@ Initial concentrations of H, and I, ase; [H₂] = 0.380 mol -0.380M [I] = 0.380 mol = 0.380 M IL IL H₂ cg) + Izeg) - 2H I cq)of Equilibrium concentrations reactants and products are; [H₂] = 0.380-0.2996 = 0.0804M (1,] - 0.380 -0.2996 = 0.0804M [HI] =